OpenStudy (anonymous):

a population grows according to p(t)=10e^t-e^(3t/2) for t>=0 a. at what rate is the population changing at t=2 and t=4 b. When is the population maximuma dn what is the maximum population c. What is the initial population? When is the population zero?

OpenStudy (anonymous):

Since the population is modeled by \(p(t)\), the rate of change of the population is given by \(p'(t)\). \[p(t)=10e^t-e^{3t/2}\\ p'(t)=10e^t-\frac{3}{2}e^{3t/2}\] For the first question, find \(p'(2)\) and \(p'(4)\). For the second, use the first derivative test to find the maxima (if any) of \(p(t)\). For the last, the initial population is the population at time \(t=0\), so find \(p(0)\). The population is zero when \(p(t)=0\), so find the \(t\) that makes this happen (if it does).

OpenStudy (anonymous):

I am unclear on the second part? I did the first derivitave test. Am i finding the local maxs? Also how do I find when adn what they are?

OpenStudy (anonymous):

Local max's occur between intervals on which \(p'(t)\) is positive and negative. So, for example, the function \(y=\dfrac{1}{1+x^2}\) has a max at \(x=0\) because its derivative is positive (and thus \(y\) is increasing) over \((-\infty,0)\), and negative (and thus decreasing) over \((0,\infty)\).